Please help me solve this if 24.3-x+17.03=31.62 find the value of x.
Answer by rothauserc(4718)

24.3 -x +17.03 = 31.62
41.33 - x = 31.62
-x = -9.71
x = 9.71
*************************************
check the answer
24.3 - 9.71 + 17.03 = 31.62
14.59 + 17.03 = 31.62
31.62 = 31.62
our answer checks